# DialSim: A Real-Time Simulator for Evaluating Long-Term Dialogue Understanding of Conversational Agents

### v1

@@ Abstract
Recent advancements in Large Language Models (LLMs) have significantly enhanced the capabilities of conversational agents, making them applicable to various fields (e.g., education). Despite their progress, the evaluation of the agents often overlooks the complexities of real-world conversations, such as real-time interactions, multi-party dialogues, and extended contextual dependencies. To bridge this gap, we introduce DialSim, a real-time dialogue simulator. In this simulator, an agent is assigned the role of a character from popular TV shows, requiring it to respond to spontaneous questions using past dialogue information and to distinguish between known and unknown information. Key features of DialSim include evaluating the agent’s ability to respond within a reasonable time limit, handling long-term multi-party dialogues, and managing adversarial settings (e.g., swap character names) to challenge the agent’s reliance on pre-trained knowledge. We utilized this simulator to evaluate the latest conversational agents and analyze their limitations. Our experiments highlight both the strengths and weaknesses of these agents, providing valuable insights for future improvements in the field of conversational AI. DialSim is available at https://github.com/jiho283/Simulator.
@@ 1 Introduction
Recent advancements in Natural Language Generation (NLG) within Large Language Models (LLMs) have significantly enhanced the capabilities of conversational agents. These agents are now integral to various fields, including customer service [behera2024cognitive, Liu2023ConversationalRS, Zhang2023UserCentricCR] and education [ait2023impact, waisberg2024large], providing personalized interactions that cater to individual preferences and interests. To ensure high user satisfaction and continuously improve these agents, it is essential to accurately evaluate their performance.
Traditionally, evaluating conversational agents has involved qualitative assessments of their responses. This process typically involves human evaluators or an LLM judging the quality of an agent’s utterances [adiwardana2020towards, kim2024prometheus, lee-etal-2023-prompted, roller2021recipes, shuster2022blenderbot, zhang-etal-2020-dialogpt] or comparing the responses between different agents (e.g., Chatbot Arena [chiang2024chatbot]). However, for conversational agents to truly engage with humans, they need to cover a wide range of communication skills, not just the ability to generate natural responses. Real-world conversations happen in real-time, often involving multiple participants and recalling previous interactions. This complexity necessitates a novel evaluation method that accurately assesses an agent’s ability to respond within a reasonable time limit, its thorough understanding of the multi-party dialogue context, and its capacity to reason across previous sessions.
Figure 1: An overall process of DialSim. Gray speech bubbles indicate predetermined utterances from the script, and white speech bubbles indicate spontaneous questions asked during the simulation. Colored speech bubbles indicate the agent’s responses to the questions. (Left) An unanswerable question. (Center) A question that references a specific time. (Right) A multi-hop question that requires understanding past sessions (i.e., the Left and Center boxes).
To address this, we propose DialSim, a dialogue simulator for real-time evaluation of a conversational agent’s long-term dialogue understanding. As illustrated in Figure 1, DialSim places an agent in the role of the main character of a TV show, engaging it in an extensive conversation based on the show’s scripted content. During this conversation, the agent is evaluated for its ability to respond appropriately to spontaneous questions from other speakers within a predefined time limit (e.g., 1s / 3s / 5s). The agent should be able to answer questions based on the information from the past dialogue and acknowledge when it does not know the answer.
@@
  • Real-Time Simulator: For a conversational agent to function effectively in real-time, the agent should be capable of updating its memory and generating responses on the fly. To evaluate this, DialSim measures the accuracy of responses within a predefined time limit. To the best of our knowledge, this is the first work that evaluates the performance of a conversational agent in a time-constrained environment.
  • Evaluating Long-Term Multi-Party Dialogue Understanding: DialSim includes multi-party dialogues with an average length of 350k tokens, making it the longest among existing long-term dialogue datasets. Throughout the simulation, the agent faces complex questions that demand comprehension and reasoning across several multi-party sessions, ensuring a thorough assessment of its understanding capabilities.
  • Randomized Test: In our simulator, there is an average of 1,300 sessions of conversations occurring over a period of 5 years as assumed in the corresponding TV show. For each session, a randomly picked character asks a question that is randomly sampled from an average of 1,000 candidates, at a random time. This setup allows us to test whether an agent can respond appropriately in an unpredictable environment, resembling real-world settings. Additionally, since a random set of questions is given for each test, repeating the tests multiple times allows us to measure the agent’s performance variability and reliability.
  • Adversarial Test: LLM-based agents are likely to have prior knowledge about these TV shows from the pre-training process (see Appendix LABEL:app:prior). Since such agents can provide answers without referencing the actual dialogue history, it is crucial to ensure that the agent relies strictly on the history for its responses. To achieve this, we developed adversarial tests that modify character names in two specific ways: by swapping their names with each other (e.g., Joey $\leftrightarrow$ Monica) or by assigning new names to them (e.g., Joey $\rightarrow$ John). This approach ensures that the agent’s responses depend strictly on the context of the dialogue history.
@@ 2 Related Works
Long-Term Dialogue Datasets    Widely used dialogue datasets include DailyDialog [li2017dailydialog], which features everyday conversations; PersonaChat [zhang2018personalizing], which contains the persona of each character; Empathetic Dialogues [rashkin2019towards], designed for emotional conversations. However, conversational agents built with these datasets were limited to single-session interactions. To address this, Multi Session Chat [xu-etal-2022-beyond] was created, featuring up to five sessions per dialogue. Despite this improvement, generating longer dialogues via crowdsourcing remained challenging. To overcome this challenge, Conversation Chronicles [jang-etal-2023-conversation] was developed by leveraging an LLM. Furthermore, LoCoMo [maharana2024evaluating] evaluates dialogue comprehension of an agent through various tasks (e.g., event summarization) in long-term dialogues. In contrast to other datasets generated through crowdsourcing or LLMs, the dialogues in DialSim are derived from TV show scripts, offering a unique set of benefits. These scripts feature multiple characters engaged in extended dialogues spanning several years, with character details (e.g., relationships) evolving over time. Leveraging these characteristics, we developed a dialogue simulator that replicates extremely long-term dialogues with evolving relationships. Table LABEL:tab:conversation_datasets provides a detailed comparison between DialSim and other long-term dialogue datasets.

### v2

# DialSim: A Real-Time Simulator for Evaluating Long-Term Multi-Party Dialogue Understanding of Conversational Agents

@@ Abstract
Recent advancements in Large Language Models (LLMs) have significantly enhanced the capabilities of conversational agents, making them applicable to various fields (e.g., education). Despite their progress, the evaluation of the agents often overlooks the complexities of real-world conversations, such as real-time interactions, multi-party dialogues, and extended contextual dependencies. To bridge this gap, we introduce DialSim, a real-time dialogue simulator. In this simulator, an agent is assigned the role of a character from popular TV shows, requiring it to respond to spontaneous questions using past dialogue information and to distinguish between known and unknown information. Key features of DialSim include evaluating the agent’s ability to respond within a reasonable time limit, handling long-term multi-party dialogues, and testing the agent’s performance under randomized questioning with a diverse and high-quality question-answer dataset. We utilized this simulator to evaluate the latest conversational agents and analyze their limitations. Our experiments highlight both the strengths and weaknesses of these agents, providing valuable insights for future improvements in the field of conversational AI. DialSim is available at https://dialsim.github.io/.
@@ 1 Introduction
Recent advancements in Natural Language Generation (NLG) within Large Language Models (LLMs) have significantly enhanced the capabilities of conversational agents. These agents are now integral to various fields, including entertainment (zhou2023characterglm; chen2024roleinteract) and education (ait2023impact; waisberg2024large), providing personalized interactions that cater to individual preferences and interests. As these agents continue to evolve and become more widely adopted, it is crucial to rigorously assess their performance in real-world scenarios to ensure they meet user expectations and function effectively.
Traditionally, the evaluation of conversational agents has relied on qualitative assessments of their responses. This process typically involves human evaluators or LLMs judging the quality of an agent’s utterances (adiwardana2020towards; zhang-etal-2020-dialogpt; roller2021recipes; shuster2022blenderbot; lee-etal-2023-prompted; kim2024prometheus) or comparing responses between different agents on platforms like Chatbot Arena (chiang2024chatbot). While these methods provide valuable insights into aspects such as naturalness and alignment with user instructions, they do not fully capture the complexities of real-world interactions.
In real-world dialogues, conversational agents encounter a range of challenges: engaging in real-time interactions, managing conversations with multiple participants, and recalling information from past dialogues. Therefore, a more comprehensive evaluation method is needed—one that accurately assesses an agent’s ability to respond within a reasonable time limit, thoroughly understand multi-party dialogue contexts, and reason across previous interactions, reflecting the complexities of real-world conversations.
Figure 1: An overall process of DialSim. Gray speech bubbles indicate predetermined utterances from the script, and white speech bubbles indicate spontaneous questions asked during the simulation. Colored speech bubbles indicate the agent’s responses to the questions. (Left) An unanswerable question. (Center) A question that references a specific time. (Right) A multi-hop question that requires understanding past sessions (i.e., the Left and Center boxes). The question is asked in the format chosen by the user, either in a multiple-choice format or as an open-ended question.
To address this, we propose DialSim, a dialogue simulator for real-time evaluation of a conversational agent’s long-term multi-party dialogue understanding. As illustrated in Figure 1, DialSim places an agent in the role of the main character of a TV show, engaging it in an extensive conversation based on the show’s scripted content. During this conversation, the agent is evaluated for its ability to respond appropriately to spontaneous questions from other speakers within a predefined time limit (e.g., 1s / 3s / 5s). The agent should be able to answer questions based on the information from the past dialogue and acknowledge when it does not know the answer.
@@
  • Time-Constrained Evaluation: For a conversational agent to function effectively in real-time, the agent must be capable of updating its memory and generating responses on the fly. To evaluate this, DialSim measures the accuracy of responses within a predefined time limit. To the best of our knowledge, this is the first work that evaluates the performance of a conversational agent in a time-constrained environment.
  • Extensive Long-Term Multi-Party Dialogue: DialSim simulates multi-party dialogues averaging 350k tokens in length, making it the longest dataset among existing long-term dialogue datasets. Throughout the simulation, the agent encounters complex questions that require comprehension and reasoning across several multi-party sessions, ensuring a thorough assessment of its long-term dialogue understanding capabilities.
  • Diverse and High-Quality Question-Answer Dataset: We generated an average of 1,000 unique questions per session using two methods. First, we collected and refined questions from a fan quiz website that covers key events in each TV show. Second, we extracted temporal knowledge graphs for each session to formulate complex, multi-hop questions based on these graphs. We employed ChatGPT-4 (openai2023gpt4) to refine the fan quizzes and extract the knowledge graphs, ensuring high quality through manual review by the authors.
  • Randomized Questioning: DialSim features an average of 1,300 conversation sessions occurring over a period of five years, as depicted in the corresponding TV show. For each session, a randomly selected character asks a question that is randomly sampled from an average of 1,000 candidates, at a random time. This setup allows us to test whether an agent can respond appropriately in a challenging and unpredictable environment, suitable for rigorous evaluation of conversational agents. Additionally, since a random set of questions is given for each test, repeating the tests multiple times allows us to measure the agent’s performance variability and reliability.
@@ 2 Related Works
Long-Term Dialogue Datasets    Widely used dialogue datasets include DailyDialog (li2017dailydialog), which features everyday conversations; PersonaChat (zhang2018personalizing), which contains the persona of each character; Empathetic Dialogues (rashkin2019towards), designed for emotional conversations. However, conversational agents built with these datasets were limited to single-session interactions. To address this, Multi Session Chat (xu-etal-2022-beyond) was created, featuring up to five sessions per dialogue. Despite this improvement, generating longer dialogues via crowdsourcing remained challenging. To overcome this challenge, Conversation Chronicles (jang-etal-2023-conversation) was developed by leveraging an LLM. Furthermore, LoCoMo (maharana2024evaluating) evaluates dialogue comprehension of an agent through various tasks (e.g., event summarization) in long-term dialogues. In contrast to other datasets generated through crowdsourcing or LLMs, the dialogues in DialSim are derived from TV show scripts, offering a unique set of benefits. These scripts feature multiple characters engaged in extended dialogues spanning several years, with character details (e.g., relationships) evolving over time. Leveraging these characteristics, we developed a dialogue simulator that replicates extremely long-term dialogues with evolving relationships. Table LABEL:tab:conversation_datasets provides a detailed comparison between DialSim and other long-term dialogue datasets.